The combustion of naphthalene (C10H8), which releases 5150.1kj/mol, is often use to calibrate calorimeters. A 1.05-g sample of naphthalene is burned in a calorimeter, producing a temperature rise of 3.86oC. Burning a 1.83-g sample of coal in the same calorimeter causes a temperature change of 4.90oC. What is the energy density of the coal?


1
Expert's answer
2021-10-22T03:50:29-0400

Molar mass of naphthalene"=10\times 12+8=128\ g"

Moles in "1.05\ g" naphthalene"=\frac{1.05}{128}"

Let the total capacity of calorimeter per unit temperature change"=C"

Using energy balance,

"\frac{1.05}{128}\times 5150.5=C\times 3.86\implies C=10.946\ KJ/\degree C"

Now coal is burnt.

Let "S" be the energy density "(" Energy stored in unit mass ")" ,

Applying energy balance,

"1.83S=10.946\times 4.90\implies S=29.31\ KJ/g"